In the parallelogram ABCD, what is the value of x?
Given: Angle A =3y Angle B= 2y-5 Angle C= 3x+3

Answer by Edwin McCravy(20060)



Two adjacent internal angles of a parallelogram are supplementary,

Therefore

     ∠A + ∠B = 180°

(3y) + (2y-5) = 180

  3y + 2y - 5 = 180

       5y - 5 = 180

           5y = 185

            y = 37

Opposite internal angles of a parallelogram have equal measure.

Therefore,

          ∠A = ∠C

          3y = 3x + 3

       3(37) = 3x + 3

         111 = 3x + 3

         108 = 3x

          36 = x 

Edwin